OpenStudy (anonymous):

If f(x) = 4x - 1 and g(x) = x2 + 2x + 2, what is f(g(x)) ?

OpenStudy (anonymous):

plug what all of g(x) into f(x) 's x

OpenStudy (anonymous):

remember that when you have f(x)= x+1 and g(x)=4x then f(g(x))=4(x+1)

OpenStudy (jit4won):

F(X) = 4x- 1 replace x with g(x) f(g(x))= 4g(X)-1 = 4(x2 + 2x + 2-1

OpenStudy (anonymous):

 plug that into x and you get \[\LARGE f(x)=4\cdot (x^2+2x+2)-1\] now solve...